# The ubiquity of variable radio emission and spin-down rates in pulsars

### v1

@@ 5.4 Transient events and interactions with planetesimals
Out of the 238 pulsars in Figure 4 with variable spin-down rates, we were able to visually identify 68 individual transient spin-down events among 26 pulsars. The absolute size and duration of these $\dot{\nu}$ events along with the inferred asteroid masses are listed in Table 4. We caution that this sample is incomplete to small transient changes in $\dot{\nu}$, which are inherently difficult to distinguish from other stochastic variations and sinusoidal changes in $\dot{\nu}$ arising from residual positional offsets. Comparing the putative asteroid masses, the smallest resulted in the second event in PSR J1833--0827 with $m\sim 2.6\times 10^{11}$ g. The third (MJD 54600-55400) event in PSR J1602--5100 is the largest with a mass of $5.32\times 10^{14}$ g. This is however smaller than the $\sim 3\times 10^{15}$ g mass inferred for the large 2005 spin-down/profile event detected in PSR J0738--4042 (Brook et al., 2014; Lower et al., 2023). Under the simplifying assumption that these objects were approximately spherical and C-type (carbonaceous) asteroids with a density of $\sim 1.57$ g cm-3 (Carry, 2012), the computed diameters range between $6.8\times 10^{3}$ to $7.5\times 10^{4}$ cm and $1.5\times 10^{6}$ cm for the 2005 event in PSR J0738--4042. Objects within this size range are exceedingly common among main-belt asteroids in the Solar System (Durda et al., 1998). If these transient spin-down events are indeed the result of pulsar-asteroid interactions, then their prevalence implies that debris discs and asteroids orbiting pulsars should be relatively common. Previous searches for planetesimals orbiting pulsars have largely been unsuccessful, though the lowest accessible mass ranges of these searches have been predominantly limited by the achieved timing precision. Surveys of 100’s of slow pulsars placed stringent limits on the presence of orbiting bodies larger than the mass of Mercury (Kerr et al., 2015; Niţu et al., 2022), while analyses of millisecond pulsar timing arrays have pushed these limits to between $\sim 0.001$-$1$ Lunar masses (Behrens et al., 2020; Niţu et al., 2024). These limits are several orders of magnitude higher than the putative asteroid masses listed in Table 4.
@@
Interactions with free-floating interstellar objects (ISOs) similar to 1I/‘Oumuamua and 2I/Borisov may present another possible origin for these events. Direct impacts between minor bodies and neutron stars have been proposed as a potential mechanism for producing both gamma-ray bursts (GRBs; Newman & Cox 1980) and fast radio bursts (FRBs; Geng & Huang 2015; Dai et al. 2016). The ISO/neutron star collision rate of $\sim 10^{7}$ Gpc-3 yr-1 (or $\sim 10^{-7}$ yr-1 for an individual neutron star) derived by Pham et al. (2024) is consistent with the observed all-sky FRB rate of $(7^{+9}_{-6})\times 10^{7}$ Gpc-3 yr-1 (Bochenek et al., 2020). Such collisions would avoid the challenges associated with forming or retaining asteroid belts around high-velocity pulsars. To test this scenario we computed the rate of observed transient spin-down events occur in our pulsar sample. For $n_{|\delta\dot{\nu}|}=70$ observed in $f_{|\delta\dot{\nu}|}=26/259$ pulsars over a total accumulated monitoring time of $T_{\rm acc}=4095.6$ yr, the rate can be derived as
@@ Acknowledgements
Murriyang, the Parkes radio telescope, is part of the Australia Telescope National Facility (https://ror.org/05qajvd42) which is funded by the Australian Government for operation as a National Facility managed by CSIRO. We acknowledge the Wiradjuri people as the traditional owners of the Observatory site. This project was supported by resources and expertise provided by CSIRO IMT Scientific Computing. Some of the data analysis behind this work was performed using the OzSTAR national HPC facility, which is funded by Swinburne University of Technology and the National Collaborative Research Infrastructure Strategy (NCRIS). MEL is supported by an Australian Research Council (ARC) Discovery Early Career Research Award DE250100508. PRB is supported by the UK Science and Technology Facilities Council (STFC), grant number ST/W000946/1. RMS acknowledges support through ARC Future Fellowship FT190100155. Part of the work was undertaken with support from the ARC Centre of Excellence for Gravitational Wave Discovery (OzGrav; CE170100004 and CE230100016). Work at NRL is supported by NASA. Pulsar research at the Jodrell Bank Centre for Astrophysics is supported by a consolidated grant from the UK STFC. This work has made use of NASA’s Astrophysics Data System. MEL is grateful to the hospitality of St Edmund Hall during visits to Oxford throughout this project.

### v2

@@ 5.4 Transient events and interactions with planetesimals
Out of the 238 pulsars in Figure 5 with variable spin-down rates, we were able to visually identify 68 individual transient spin-down events in 26 pulsars. The absolute size and duration of these $\dot{\nu}$ events along with the inferred asteroid masses are listed in Table 4. We caution that this sample is incomplete to small transient changes in $\dot{\nu}$, which are inherently difficult to distinguish from other stochastic variations and sinusoidal changes in $\dot{\nu}$ arising from residual positional offsets. Comparing the putative asteroid masses, the smallest resulted in the second event in PSR J1833--0827 with $m\sim 2.6\times 10^{11}$ g. The third (MJD 54600-55400) event in PSR J1602--5100 is the largest with a mass of $5.32\times 10^{14}$ g. This is however smaller than the $\sim 3\times 10^{15}$ g mass inferred for the large 2005 spin-down/profile event detected in PSR J0738--4042 (Brook et al., 2014; Lower et al., 2023). Under the simplifying assumption that these objects were approximately spherical and C-type (carbonaceous) asteroids with a density of $\sim 1.57$ g cm-3 (Carry, 2012), the computed diameters range between $6.8\times 10^{3}$ to $7.5\times 10^{4}$ cm and $1.5\times 10^{6}$ cm for the 2005 event in PSR J0738--4042. Objects within this size range are exceedingly common among main-belt asteroids in the Solar System (Durda et al., 1998). If these transient spin-down events are indeed the result of pulsar-asteroid interactions, then their prevalence implies that debris discs and asteroids orbiting pulsars should be relatively common. Previous searches for planetesimals orbiting pulsars have largely been unsuccessful, though the lowest accessible mass ranges of these searches have been predominantly limited by the achieved timing precision. Surveys of 100’s of slow pulsars placed stringent limits on the presence of orbiting bodies larger than the mass of Mercury (Kerr et al., 2015; Niţu et al., 2022), while analyses of millisecond pulsar timing arrays have pushed these limits to between $\sim 0.001$-$1$ Lunar masses (Behrens et al., 2020; Niţu et al., 2024). These limits are several orders of magnitude higher than the putative asteroid masses listed in Table 4.
@@
Interactions with free-floating interstellar objects (ISOs) similar to 1I/‘Oumuamua and 2I/Borisov may present another possible origin for these events. Direct impacts between minor bodies and neutron stars have been proposed as a potential mechanism for producing both gamma-ray bursts (GRBs; Newman & Cox 1980) and fast radio bursts (FRBs; Geng & Huang 2015; Dai et al. 2016). The ISO/neutron star collision rate of $\sim 10^{7}$ Gpc-3 yr-1 (or $\sim 10^{-7}$ yr-1 for an individual neutron star) derived by Pham et al. (2024) is consistent with the observed all-sky FRB rate of $(7^{+9}_{-6})\times 10^{7}$ Gpc-3 yr-1 (Bochenek et al., 2020). Such collisions would avoid the challenges associated with forming or retaining asteroid belts around high-velocity pulsars. To test this scenario we computed the rate of observed transient spin-down events occur in our pulsar sample. For $n_{|\delta\dot{\nu}|}=70$ events observed in $f_{|\delta\dot{\nu}|}=27/260$ pulsars (note, this includes the two events in PSR J0738--4042) over a total accumulated monitoring time of $T_{\rm acc}=4095.6$ yr, the rate can be derived as
@@ Acknowledgements
Murriyang, the Parkes radio telescope, is part of the Australia Telescope National Facility (https://ror.org/05qajvd42) which is funded by the Australian Government for operation as a National Facility managed by CSIRO. We acknowledge the Wiradjuri people as the traditional owners of the Observatory site. This project was supported by resources and expertise provided by CSIRO IMT Scientific Computing. Some of the data analysis behind this work was performed using the OzSTAR national HPC facility, which is funded by Swinburne University of Technology and the National Collaborative Research Infrastructure Strategy (NCRIS). MEL is supported by an Australian Research Council (ARC) Discovery Early Career Research Award DE250100508. PRB is supported by the UK Science and Technology Facilities Council (STFC), grant number ST/W000946/1. RMS acknowledges support through ARC Future Fellowship FT190100155. Part of the work was undertaken with support from the ARC Centre of Excellence for Gravitational Wave Discovery (OzGrav; CE170100004 and CE230100016). Work at NRL is supported by NASA. Pulsar research at the Jodrell Bank Centre for Astrophysics is supported by a consolidated grant from the UK STFC. This work has made use of NASA’s Astrophysics Data System, and the following open-source software packages that were not cited elsewhere: matplotlib (Hunter, 2007), numpy (Harris et al., 2020) and scipy (Virtanen et al., 2020). MEL is grateful to the hospitality of St Edmund Hall during visits to Oxford throughout this project.